OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root of quantity x divided by nine. - 4

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

\(\Large f(x)=\sqrt[3]{ \frac x9}-4\) make f(x) = y \(\Large y=\sqrt[3]{ \frac x9}-4\) swap x and y terms \(\huge \color{red}x=\sqrt[3]{ \frac {\color{red}y}9}-4\) now put x in terms of y \(\Large x=\sqrt[3]{ \frac y9}-4\) \(\Large x+4=\sqrt[3]{ \frac y9}\) \(\Large (x+4)^3=\frac y9\) \(\Large 9(x+4)^3=y\) \(\Large 9(x+4)^3=f^{-1}(x)\) \(\Large f^{-1}(x) = 9(x+4)^3\) does this make sense @PHUNISH ?

OpenStudy (anonymous):

Ohh yeah for some reason I kept cubing the x idk how or why thanks though!

OpenStudy (jack1):

np dude ;)

OpenStudy (anonymous):

I have a couple more I got wrong if you could explain them to me, ill tag you in them

OpenStudy (jack1):

cools ;)